Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Permutaion

Expert replies
by SwatiDenre » Fri Jun 10, 2011 12:50 pm
Let A, B and C denotes separate generic traits. Suppose the probability that an individual will exhibit trait A is 1/2 and the probability that he exhibits B trait is 3/4 and the probability that he exhibits trait C is 3/5. What is the probability that he exhibits only one of the traits?

A. 1/4
B. 11/40
C. 1/6
D. 17/36
E. 11/36
Join the discussion
Source: — Problem Solving |

by jainnikhil02 » Fri Jun 10, 2011 1:10 pm
IMO A

Let me know if i am wrong
Nikhil K Jain
____________________

"Life is all about timing" Don't waste your and others time.
Join the discussion

by SwatiDenre » Fri Jun 10, 2011 1:46 pm
Ans - B

But how did u get through it.
Join the discussion

by pemdas » Fri Jun 10, 2011 2:28 pm
this one is mixed concept q., probability+permutation

P(not A)=1-1/2=1/2
P(not B)=1-3/4=1/4
P(not C)=1-3/5=2/5

P(only A)=P(A)*P(not B)*P(not C)=1/2*1/4*2/5=1/20
P(only B)=P(B)*P(not A)*P(not C)=3/4*1/2*2/5=3/20
P(only C)=P(C)*P(not A)*P(not B)=3/5*1/2*1/4=3/40

1/20+3/20+3/40=11/40
SwatiDenre wrote:Let A, B and C denotes separate generic traits. Suppose the probability that an individual will exhibit trait A is 1/2 and the probability that he exhibits B trait is 3/4 and the probability that he exhibits trait C is 3/5. What is the probability that he exhibits only one of the traits?

A. 1/4
B. 11/40
C. 1/6
D. 17/36
E. 11/36
Success doesn't come overnight!
Join the discussion

by cans » Fri Jun 10, 2011 8:11 pm
P(A)=1/2
p(B)=3/4
p(C)=3/5
p(one of A,B,C) = p(A & not B * not C) + p(B & not A * not C) + p(C & not B * not A)
= (1/2)(1/4)(2/5) + (3/4)(1/2)(2/5) + (3/5)(1/4)(1/2)
=11/40
If my post helped you- let me know by pushing the thanks button ;)

Contact me about long distance tutoring!
[email protected]

Cans!!
Join the discussion

by jainnikhil02 » Fri Jun 10, 2011 8:56 pm
approch was rite but done calculation mistakte.. :(
Nikhil K Jain
____________________

"Life is all about timing" Don't waste your and others time.
Join the discussion